OpenStudy (anonymous):

Which equation represents the line passing through the points (2, −2) and (−5, 12)? 2x − y = 2 2x + y = 2 x + 2y = 2 x − 2y = 2

OpenStudy (anonymous):

How do we find the equation of a line? Find the slope first, introduce x and y. \[(y_2-y_1)/(x_2-x_1)\] \[(y+2)=slope \times (x-2)\]

OpenStudy (anonymous):

ya im still lost

OpenStudy (anonymous):

Joel, find the slope, then rearrange by keeping the x and y on one side.

OpenStudy (anonymous):

\[slope = (12-(-2))/(-5-2)\] Now your equation will be the answer you get above times (x-2)=(y+2)

OpenStudy (anonymous):

if you're still lost try just plugging in those points into each equation and see which one holds true...

OpenStudy (anonymous):

look ive been doing algibra since 10am this mornig and its 8:33pm ive pretty much fried my brain so im trying to finish the course im on the last thing so im going to take a break or im going to break